1)3x + 5 - 5(x+1)=6x +7
2. 2(x - 8)=3x - 26
3. 4(x+7)=2(x+12)+2(x+1)

Answer by ikleyn(52776)

.


            In this post,  I will solve equation  2  for you,  with full explanations.


Your starting equation is

    2(x-8) = 3x - 26


Simplify it step by step

    2x - 16 = 3x - 26


Collect the terms with x on the right side.
Collect all other terms (constant terms) in the left side.

By moving the terms over the equation sign, from left side to right side and vice versa,
change the sign of the term to the opposite one, each time.

    26 - 16 = 3x - 2x


Combine and compute like terms

      10   =    x.


ANSWER.  x = 10.

Solved.

----------------

On solving single linear equation in one unknown see the lesson
    - HOW TO solve a linear equation
in this site.

Consider this lesson as your textbook,  handbook,  guide,  tutorials and  (free of charge)  home teacher.

Happy learning (!)


/////////////////


To get clear explanations,  it is  ALWAYS  better to post your problems  SEPARATELY
and do not pack more than one problem per post.


It is the policy of this web-site.


It guarantees the best service for you  ( ! )


\\\\\\\\\\\\\\\


About the rules,  see this web-page

https://www.algebra.com/tutors/students/ask.mpl?action=ask_question&topic=Equations&return_url=http://www.algebra.com/algebra/homework/equations/

from which you post your problems.